5263=235,000 / 100 - x
140,000 = 4,250,00 / x-50

Answer by stanbon(75887)

5263=235,000 / 100 - x
100 - x = 235000/5263
x = -44.65+100
x= 55.35
==================

140,000 = 4,250,00 / x-50
x-50 = 4250000/140000
x-50 = 30.36
x = 80.36
==================
Cheers,
Stan H.